question1:

What is the magnetic field in the axis through the center of a square current loop (width L) as a function of the height (y) above the plane of the loop?


Question 2

What is the magnetic field due to a current segment at a target point of coordinates (x,y)? Set up the integral and the cross product and integrate. No need to simplify the answer.


Question 1 What is the magnetic Seld in the axis through the center of a square current loop (width L) as a function of the height () above the plane of the loop? target point height, 2 length & wicm L current Question 2 What is the magneic field due to a current segment at a target point of coordinates (xy)? Set up the integral and the cross product and integrate. No need to simplify the answer. target port(x,y) current current, I length, L page 1
Answer #1

Using\:Biot-Savat\:equation:

dB=\frac{\mu }{4\pi }\frac{I\left (d\vec{L}\times \vec{r} \right )}{r^{3}}

Let\:\theta \:and\:\alpha ,\:the\:angles\:respect\:the\:axis.

dB=\frac{\mu IdL\sin \theta }{4\pi r^{2} }=\frac{\mu IdL\sin \alpha }{4\pi r^{2}}=\frac{\mu I}{4\pi }\frac{dxR}{r^{3}}

Integrating:

B=\frac{\mu Ir}{4\pi }\int_{-\frac{L}{2}}^{\frac{L}{2}}\frac{dx}{\left ( x^{2}+R^{2} \right )^{\frac{3}{2}}}=\frac{\mu Ir}{4\pi }\frac{x}{R^{2}\left ( x^{2}+R^{2} \right )^{\frac{1}{2}}}\:|_{\frac{-L}{2}}^{\frac{L}{2}}

B=\frac{\mu I}{4\pi R}\left (\cos \alpha_1-\cos \alpha _2 \right )

and

R=\sqrt{y^{2}+\left (\frac{L}{2} \right )^{2}}

\cos \alpha _1-\cos \alpha _2=\frac{\frac{L}{2}}{\sqrt{R^{2}+\frac{L^{2}}{4}}}=\frac{L}{2\sqrt{y^{2}+\frac{L^{2}}{2}}}

substituting:

B=\frac{\mu I}{4\pi R}\frac{a}{\sqrt{y^{2}+\frac{L^{2}}{2}}}

This\:is\:the\:result\:for\:each\:side,\:so,\:total\:B\:is:

B = 4B sin B = 4B sin 2R